OpenStudy (anonymous):

Calculate the density of a rectangular piece of wood from the following data: Length: 25 cm. Width: 10 cm. Height: 4 cm. Mass: 15 g a. 0.015 g/cm3 c. 0.015 g/ml b. 10.015 g/ml d. None of the above

OpenStudy (anonymous):

\[ \text{Volumetric density} = \rho = \frac{\text{mass}}{\text{volume}} \ \ \propto \ g/cm^3\] \[V_{rectangle} = L \cdot W \cdot H \]